Is the given number a solution of the equation?
9k = 10 - k; -1
-7r - 15 = -2r; -3
-3p = 4p + 35; -5
8 -e = 2e - 16; 8
5 - 15s = 8 - 16s; 3
2(x - 2) - 5x = 5(2 - x); 7
6a + 3 = 3(3a - 2); 4

Answer by funmath(2933)

You determine if a number is a solution to an equation by substituting the number into the equation for the variable and seeing if both sides actually equal each other once you simplify them.
9k=10-k;-1
9(-1)=10-(-1)
-9=10+1
-9=11 No.
-----------------
7r-15=-2r;-3
7(-3)-15=-2(-3)
-21-15=6
-36=6 No
------------------
8-e=2e-16;8
8-(8)=2(8)-16
8-8=16-16
0=0 Yes
-------------------
5-15s=8-16s;3
5-15(3)=8-16(3)
5-45=8-48
-40=-40 yes
-------------------
2(x-2)-5x=5(2-x);7
2(7-2)-5(7)=5(2-7)
2(5)-5(7)=5(-5)
10-35=-25
-25=-25 Yes
--------------------
6a+3=3(3a-2);4
6(4)+3=3(3(4)-2)
6(4)+3=3(12-2)
6(4)+3=3(10)
24+3=30
27=30 N0